OpenStudy (anonymous):

solve each equation for all real values of x. ((solving trigonometric equations) 2sinx+cscx=3

OpenStudy (anonymous):

\[2\sin x +\csc x = 3\]\[2\sin x +\frac{ 1 }{ \sin x } = 3\]Multiply both sides by sin x.\[2\sin^2x +1=3\sin x \]\[2\sin^2 x - 3\sin x +1= 0\]\[\left( 2\sin x - 1 \right) \left( \sin x -1 \right)= 0\] then solve for x. hope that helps

OpenStudy (anonymous):

thank you so much!!!